OpenStudy (anonymous):

The earth rotates through 1 complete revolution every 1,440 minutes. Since the axis of rotation is perpendicular to the equator, you can think of a person standing on the edge of a disc that is rotating through 1 complete revolution every 1,440 minutes. Find the angular velocity of a person standing on the equator

OpenStudy (tkhunny):

\(\dfrac{1\;rev}{1440\;min}\cdot\dfrac{2\pi}{1\;rev}\)